Using a stadard 52 card deck (no jokers) find the probability that one card is picked from the deck, that card is either below a five or it is a red.
Answer by venugopalramana(3286)

Using a stadard 52 card deck (no jokers) find the probability that one card is picked from the deck, that card is either below a five or it is a red.
THERE ARE 26 RED CARDS.
BELOW 5 THERE ARE CARDS WITH 4,3 AND 2 DENOMINATIONS =3*4=12 CARDS (ACE IS EXCLUDED..IF IT IS CONSIDERED AS 1 THEN THAT HAS TO BE INCLUDED IN WHICH CASE THERE ARE 4*4 =16 CARDS.) . BUT THEY INCLUDE 6 (OR WITH ACE 8 ) RED CARDS,WHICH ARE COUNTED ABOVE.SO EXCLUDING THESE WE HAVE 6 ( OR 8) CARDS BELOW 5.
SO SUCCESS POSSIBILITIES =26+6(8)=32(34)
TOTAL POSSIBILITIES =52
PROBABILITY OF SUCCESS = 32/52=8/13...OR....34/52=17/26